# Feedforward in Generative AI: Opportunities for a Design Space

### v1

@@ Abstract.
Generative AI (GenAI) models have become more capable than ever at augmenting productivity and cognition across diverse contexts. However, a fundamental challenge remains as users struggle to anticipate what AI will generate. As a result, they must engage in excessive turn-taking with the AI’s feedback to clarify their intent, leading to significant cognitive load and time investment. Our goal is to advance the perspective that in order for users to seamlessly leverage the full potential of GenAI systems across various contexts, we must design GenAI systems that not only provide informative feedback but also informative feedforward—designs that tell users what AI will generate before the user submits their prompt. To spark discussion on feedforward in GenAI, we designed diverse instantiations of feedforward across four GenAI applications: conversational UIs, document editors, malleable interfaces, and automation agents, and discussed how these designs can contribute to a more rigorous investigation of a design space and a set of guidelines for feedforward in all GenAI systems.
@@
journalyear: 2025copyright: acmcopyrightconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 0000000.0000000isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interaction paradigmsccs: Human-centered computing HCI theory, concepts and models
@@ 1. Introduction
Despite its potential, users inevitably face challenges in trying to anticipate GenAI’s responses and aligning their prompts to match their needs. Subramonyam et al. have labeled this set of challenges as the gulf of envisioning, which describes the user’s gap in identifying and articulating their desired task, providing effective instructions, and anticipating the model’s response (Subramonyam et al., 2024). This gap makes it difficult for novices to formulate effective prompts to AI, understand the cause of misalignment, and persist in iterating and refining their prompts (Zamfirescu-Pereira et al., 2023).
@@
While researchers have explored various approaches to help users better learn and anticipate AI’s responses by providing dedicated spaces for learning and experimenting with prompts (Arawjo et al., 2024; Angert et al., 2023), generating multiple output variations in a single space (Suh et al., 2024; Gero et al., 2024), asking clarifying questions (OpenAI, 2025b), and reducing friction in modifying generated outputs through direct manipulation (Vaithilingam et al., 2024b; Masson et al., 2024b; Jiang et al., 2023), these systems primarily focus on offering feedback—helping users understand what happened after they submit a prompt. While feedback helps users develop an understanding of the GenAI system through exploration and experience, users cannot always anticipate the outcome of their prompts and must rely on iterative trial-and-error. This process can incur significant cognitive load and time investment, especially for heavy tasks like generating user interfaces or automating web tasks. If we want users to easily leverage the potential of GenAI systems across diverse contexts seamlessly, they must be able to anticipate AI outputs without solely relying on prior experience and feedback, and instead tune their response in faster increments.
@@
To achieve this, we aim to advance the perspective that GenAI must not only provide informative feedback but also informative feedforward—designs that help users anticipate outcomes before performing an action (Vermeulen et al., 2013). Providing such feedforward in GenAI systems can help users understand what to expect from the AI model before submitting their prompt. This can reduce excessive conversational exchanges, bridge the abstraction gap between users and AI (Liu et al., 2023; Subramonyam et al., 2024; Vaithilingam et al., 2024a), and encourage meta-cognition when constructing prompts (Tankelevitch et al., 2024). Reflecting this need, the HCI community has begun exploring feedforward designs for GenAI, examining trade-offs between the amounts of detail provided and cognitive load when assessing generated content (Zhu-Tian et al., 2024; Dang et al., 2022b). However, fully understanding these trade-offs across diverse modalities and contexts requires rigorous investigation, as the community must systematically study feedforward designs in diverse GenAI systems, assess user benefits and challenges across design dimensions, and establish comprehensive design guidelines.
@@
This paper aims to spark discussion on how we may achieve these goals, enabling users to effectively engage and interact with all forms of GenAI systems. To lay the groundwork for this discussion, we designed diverse instantiations of feedforward across four GenAI applications: conversational UIs, document editors, malleable interfaces, and automation agents. We then discussed the designs and implementations needed to further develop a comprehensive design space for informative feedforward GenAI systems.
@@
We invite researchers in this community to critically examine past and present GenAI interfaces through the lens of feedforward, draw from feedforward literature in other domains (Bau and Mackay, 2008; Coppers et al., 2019; Terry and Mynatt, 2002; Muresan et al., 2023), explore new interaction techniques and frameworks for delivering informative feedforward, and work toward concrete design and implementation strategies applicable to all GenAI systems.
2. Feedforward for Generative AI
@@
To explore various ways to design feedforward in GenAI systems, we first detail our design of feedforward in an LLM-powered conversational interface. We then describe scenarios for three other designs of feedforward in GenAI applications: document editors, malleable interfaces, and agent automation systems.
@@ 2.1.1. Feedforward Components
Research has shown that users benefit from multiple representations when interacting with LLMs (Jiang et al., 2023; Suh et al., 2023). Therefore, we developed feedforward components that each aim to provide a specific representation that would help users anticipate the LLM’s response. As the user types a prompt and pauses for a brief moment, the conversational UI generates feedforward content inside two feedforward components. The first component presents key topics and a high-level outline of anticipated response (Fig. 1.1a), while the second component presents a minimap view presenting paragraphs in the form of blocks (Fig. 1.1b). These representations help users anticipate the topics and subtopics the LLM will generate, as well as whether its response might be too short or too long.
@@ 2.1.2. Manipulating Feedforward Content
We enable users to manipulate feedforward content in two ways. First, when users edit their prompt, the feedforward content updates to reflect the changes. This allows users to identify misalignments between their intent and the feedforward content, enabling them to add details before submitting their final prompt. For instance, they can clarify an ambiguous term (Fig. 1.2) or request a shorter response to reduce information overload (Fig. 1.3). Such instant updates to the feedforward components can spark reflection, such as realizing that it may be useful to understand both the advantages and limitations of Wizard of Oz Prototyping rather than just its definition.
@@ 2.2. Application Scenarios
Our goal is to explore opportunities for designing feedforward in all GenAI interfaces—beyond conversational UIs. Thus, we present three additional application scenarios that showcase various feedforward designs across diverse contexts.
@@ 3. Future Work
Our goal is to establish feedforward as a fundamental design component in all GenAI systems. While this paper presents several feedforward designs, further exploration is needed to effectively design and implement feedforward across all GenAI applications.
In order to achieve this, we must streamline the design process of feedforward for GenAI by establishing a comprehensive design space. From our prototype explorations across four applications—conversational UIs, document editors, malleable interfaces, and agent automations—we recognize three potential design dimensions: representation, level of detail, and manipulability. First, we designed feedforward representations in the form of outlines, minimaps, lists of operations, example phrases, and multiple cursors. We anticipate that a more rigorous investigation of feedforward will reveal additional representations suited for various contexts, such as data visualizations, UI wireframes, and diagrams. Second, we explored different ways to present varying levels of detail in feedforward. For example, the conversational UI displays an outline summarizing key topics, while the minimap organizes topics into paragraphs. A rigorous investigation on the optimal levels of detail for various feedforward representations can further help users anticipate AI responses without unnecessary cognitive burden. Lastly, we explored how users can manipulate feedforward content either by revising their prompts or directly resizing, repositioning, or selecting elements. We aim to explore additional interaction techniques that could enhance feedforward designs. Future work should build on this preliminary design space by gathering, analyzing, and critiquing a broader collection of GenAI systems to identify potential variations of feedforward across more contexts.

### v2

@@ Abstract.
Generative AI (GenAI) models have become more capable than ever at augmenting productivity and cognition across diverse contexts. However, a fundamental challenge remains as users struggle to anticipate what AI will generate. As a result, they must engage in excessive turn-taking with the AI’s feedback to clarify their intent, leading to significant cognitive load and time investment. Our goal is to advance the perspective that in order for users to seamlessly leverage the full potential of GenAI systems across various contexts, we must design GenAI systems that not only provide informative feedback but also informative feedforward—designs that tell users what AI will generate before the user submits their prompt. To spark discussion on feedforward in GenAI, we designed diverse instantiations of feedforward across four GenAI applications: conversational UIs, document editors, malleable interfaces, and agent automations, and discussed how these designs can contribute to a more rigorous investigation of a design space and a set of guidelines for feedforward in all GenAI systems.
@@
journalyear: 2025copyright: acmcopyrightconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japanisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interaction paradigmsccs: Human-centered computing HCI theory, concepts and models
@@ 1. Introduction
Despite its potential, users inevitably face challenges in trying to anticipate GenAI’s responses and aligning their prompts to match their needs. Subramonyam et al. have labeled this set of challenges as the gulf of envisioning, which describes the user’s gap in identifying and articulating their desired task, providing effective instructions, and anticipating the model’s response (Subramonyam et al., 2024). This gap makes it difficult for novices to formulate effective prompts to AI and understand the cause of misalignment (Zamfirescu-Pereira et al., 2023).
@@
While researchers have explored various approaches to help users better learn and anticipate AI’s responses with dedicated spaces for experimenting with prompts (Arawjo et al., 2024; Angert et al., 2023), generating multiple outputs in a single space (Suh et al., 2024; Gero et al., 2024), asking clarifying questions (OpenAI, 2025b), and providing widgets to modify generated outputs post-hoc (Vaithilingam et al., 2024b; Masson et al., 2024b; Jiang et al., 2023), these systems primarily focus on offering feedback—helping users understand what happened after they submit a prompt. While feedback helps users develop an understanding of the system through exploration, users cannot always anticipate the outcome of their prompts. Instead, they must rely on iterative trial-and-error, which can incur significant cognitive load and time investment. If we want users to easily leverage the potential of GenAI systems across diverse contexts, they must be able to anticipate AI outputs without solely relying on feedback.
@@
To achieve this, we aim to advance the perspective that GenAI must not only provide informative feedback but also informative feedforward—designs that help users anticipate outcomes before performing an action (Vermeulen et al., 2013). Providing such feedforward in GenAI systems can help users understand what to expect from the AI model before submitting their prompt. This can reduce excessive conversational exchanges, bridge the abstraction gap between users and AI (Liu et al., 2023; Subramonyam et al., 2024; Vaithilingam et al., 2024a), and encourage meta-cognition when constructing prompts (Tankelevitch et al., 2024).
The HCI community has extensively explored feedforward design in various domains, such as gestures (Bau and Mackay, 2008), VR/AR (Muresan et al., 2023), widgets (Coppers et al., 2019; Terry and Mynatt, 2002), and proactive systems (Allen et al., 1999). Furthermore, the community has begun exploring feedforward designs for GenAI, examining trade-offs between the level of detail provided and the cognitive load involved in assessing generated code and images (Zhu-Tian et al., 2024; Dang et al., 2022b). However, GenAI’s expanding capabilities are shaping an increasingly broad design space of GenAI interfaces. This rapid growth calls for a more systematic approach to designing effective feedforward across diverse domains—uncovering opportunities for novel interaction techniques, assessing user benefits and challenges across design dimensions, and establishing comprehensive design guidelines.
@@
This paper aims to spark discussion on how we may achieve these goals, enabling users to effectively engage and interact with all forms of GenAI systems. To lay the groundwork for this discussion, we designed diverse instantiations of feedforward across four GenAI applications: conversational UIs, document editors, malleable interfaces, and agent automations. We then discussed the designs and implementations needed to further develop a comprehensive design space for informative feedforward GenAI systems.
@@
We invite researchers in this community to critically examine past and present GenAI interfaces through the lens of feedforward, draw from feedforward literature in other domains, explore new interaction techniques and frameworks for delivering informative feedforward, and work toward concrete design and implementation strategies applicable to all GenAI systems.
2. Feedforward in Generative AI
@@
To explore various ways to design feedforward in GenAI systems, we first explore feedforward in an LLM-powered conversational interface. We then describe scenarios for three other designs of feedforward in GenAI applications.
@@ 2.1.1. Feedforward Components
Research has shown that users benefit from multiple representations when interacting with LLMs (Jiang et al., 2023; Suh et al., 2023). Therefore, we developed feedforward components that each aim to provide a specific representation that would help users anticipate the LLM’s response. As the user types a prompt, the conversational UI generates feedforward content inside two feedforward components: 1) a key topic phrase alongside a high-level outline of the anticipated response (Fig. 1.1a) and 2) a minimap view with paragraph blocks (Fig. 1.1b). These representations help users anticipate the topics and subtopics the LLM will generate, as well as whether its response might be too short or too long.
@@ 2.1.2. Manipulating Feedforward Content
We enable users to manipulate feedforward content in two ways. First, when users edit their prompt, the feedforward content updates to reflect the changes. This allows users to identify misalignments between their intent and the feedforward content, encouraging them to add details to the prompt before submitting it. For instance, they can clarify an ambiguous term (Fig. 1.2) or request a shorter response (Fig. 1.3). These instant updates to the feedforward components may also spark reflection, such as realizing that it may be useful to understand the advantages and limitations of Wizard of Oz Prototyping rather than just its definition.
@@ 2.2. Application Scenarios
Our goal is to explore opportunities for designing feedforward in all GenAI interfaces—beyond conversational UIs. In this paper, we explore three applications to showcase various feedforward designs across diverse contexts: document editors, malleable interfaces, and agent automations.
@@ 3. Future Work
Our goal is to establish feedforward as a fundamental design component in all GenAI systems. While this paper presents several feedforward designs, we envision a more comprehensive design space to represent and guide feedforward design across all GenAI applications. Based on our four prototypes, we identify three potential design dimensions for GenAI feedforward: representation, level of detail, and manipulability.
First, our examples implemented feedforward representations in the form of outlines, minimaps, lists of operations, example phrases, and multiple cursors. We may identify categories of representations that help distinguish which feedforward representations are more useful for different use cases. For instance, while a list of operations might inform users about the type of UI a GenAI system will generate, a wireframe might better communicate the structure and layout of the UI.
Second, we explored different ways to present varying levels of detail in feedforward. For example, the conversational UI displays an outline summarizing key topics, while the minimap omits textual details and instead presents blocks of paragraphs. This dimension aligns with previous research on feedforward (Bau and Mackay, 2008). A more rigorous investigation could explore optimal levels of detail for different types of feedforward representations in GenAI, as well as allowing users to define the level and type of detail themselves.
Lastly, we explored how users can manipulate feedforward content, either by revising their prompts or by directly resizing, repositioning, or selecting elements. We aim to investigate additional interaction techniques that enhance user control and engagement with feedforward designs.
Future work should expand on this preliminary design space by gathering, analyzing, and critiquing a broader range of GenAI systems to identify variations of feedforward across diverse contexts.